# MoLAN: A Unified Modality-Aware Noise Dynamic Editing Framework for Multimodal Sentiment Analysis

### v1

@@ Abstract
Multimodal Sentiment Analysis aims to integrate information from various modalities, such as audio, visual, and text, to make complementary predictions. However, it often struggles with irrelevant or misleading visual and auditory information. Most existing approaches typically treat the entire modality information (e.g., a whole image, audio segment, or text paragraph) as an independent unit for feature enhancement or denoising. They often suppress the redundant and noise information at the risk of losing critical information. To address this challenge, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. Specifically, MoLAN performs modality-aware blocking by dividing the features of each modality into multiple blocks. Each block is then dynamically assigned a distinct denoising strength based on its noise level and semantic relevance, enabling fine-grained noise suppression while preserving essential multimodal information. Notably, MoLAN is a unified and flexible framework that can be seamlessly integrated into a wide range of multimodal models. Building upon this framework, we further introduce MoLAN+, a new multimodal sentiment analysis approach. Experiments across five models and four datasets demonstrate the broad effectiveness of the MoLAN framework. Extensive evaluations show that MoLAN+ achieves the state-of-the-art performance. The code is publicly available at https://github.com/betterfly123/MoLAN-Framework.
Introduction
The development demands of Artificial General Intelligence increase attention toward Multimodal Sentiment Analysis (MSA). MSA aims to integrate information from various modalities, such as text, visual, and audio data, to achieve a more comprehensive and accurate understanding of the emotions expressed by users or the environment (Zadeh et al. 2018b; Morency, Mihalcea, and Doshi 2011). MSA holds significant academic value in advancing multimodal learning and cross-modal understanding, and offers broad industrial applications in areas such as human-computer interaction and mental health monitoring (Das and Singh 2023; Pantic and Rothkrantz 2003; Zhou et al. 2015).
@@ Introduction
Most existing methods for MSA achieve impressive improvements by leveraging the synergistic effects of multiple modalities, for example, text, audio, and visual data (Han, Chen, and Poria 2021; Wang 2022; Mai, Zeng, and Hu 2022; Wu et al. 2024; Xie et al. 2024; Li and Liu 2025; Lin et al. 2022). However, due to irrelevant and misleading noise, the application of MSA systems in real-world scenarios remains challenging. It has been demonstrated that misleading noise information can seriously interfere with the model’s representation learning and further lead to performance degradation (Li and Li 2025; Liu 2024; Mao et al. 2023). To deal with multiple noise patterns, a straightforward simple solution is to train individual models from scratch for each type of noise (Yuan et al. 2021; Li et al. 2020). This strategy relies on prior knowledge of characteristic noise, which is typically unavailable in practical applications. An intuitive improvement is to design a unified model to adaptively perceive noise features (Zeng, Zhou, and Liu 2022b). However, there are significant differences among various types of modality-specific noise. For instance, noise in the visual modality is usually caused by non-target person interference, while noise in the audio modality mainly comes from the environmental background. Consequently, a model optimized for a specific type of feature noise may yield suboptimal performance when applied to other noise modalities. To address this limitation, researchers design different denoising methods based on the characteristics of different noise for each noise pattern. (Yuan et al. 2023, 2024). Notably, the intensity and contribution of noise differ across regions within each modality. Therefore, those approaches that denoise the entire modality as a whole may suppress noise information at the cost of losing essential information.
@@
As shown in Figure 1, different modalities contain noise that is inconsistent with the sentiment labels, and the intensity of this noise varies across different regions. Specifically, the visual modality contains visual noise that contradicts Ross’s emotional state, such as people’s smiles in the background, which constitute strong noise. While other parts of the background represent weak noise. Similarly, the audio contains segments of strong noise, such as laughter, that conflict with the sentiment label. The inconsistency of the noise distribution highlights the importance of fine-grained denoising that can distinguish between varying levels of noise intensity and selectively suppress noise information without losing essential information. Therefore, the challenge of this paper is how to perform fine-grained noise dynamic editing on different modalities, so as to remove noise information while retaining information that is beneficial to MSA.
@@
To address these issues, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. To achieve dynamic fine-grained denoising, MoLAN employs a blocking strategy that divides each modality into different sub-blocks. In each block, the denoising strength is dynamically computed based on the noise level of the block. This approach allows the model to apply varying degrees of denoising across different blocks, thereby enhancing its ability of selective noise editing while preserving essential information in each modality. In addition, due to the heterogeneity between different modalities (Fan et al. 2024; Wei et al. 2023), we adopt different block strategies for different modalities. The visual modality uses two-dimensional blocks, while the audio modality uses one-dimensional blocks. Following previous works (Li and Li 2025; Zhang et al. 2023; Lin and Hu 2022), we choose text as the main modality to calculate the denoising strength. Additionally, MoLAN is a unified framework that can be flexibly integrated into different MSA methods to improve the upper limit of model performance. Based on the MoLAN, we propose MoLAN+, which uses denoised information to update the cross-attention between modalities and guide the model to focus on important information for MSA. MoLAN+ further introduces denoising-driven contrastive learning to encourage the model to generate higher quality features, improving the performance of MSA task. The key contributions of our work are as follows:
@@
  • To address the noise problem, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. MoLAN performs modality-aware blocking by dividing the modality features into multiple blocks. Each block is dynamically assigned a distinct denoising strength, enabling fine-grained noise editing. Additionally, MoLAN can be flexibly integrated into various MSA models.
@@
  • We conduct experiments on five models and four datasets to demonstrate the broad effectiveness of the MoLAN framework. Additionally, extensive evaluations on four benchmark multimodal datasets show that MoLAN+ achieves the state-of-the-art performance.
@@
Figure 2: An illustration of our proposed MoLAN framework and MoLAN+ method. The purple box above represents the MoLAN framework, and the below represents the entire process of the framework-based MoLAN+ method. The MoLAN framework shown above provides a detailed description of the MoLAN block presented in MoLAN+ method.
Related Work
Multimodal Sentiment Analysis (MSA)
MSA aims to enable machines to understand human emotions by leveraging multimodal signals. Early MSA research primarily employs methods such as TFN (Zadeh et al. 2017) and LMF (Liu and Shen 2018) to obtain multimodal fused representations through different feature extraction networks. As technology advances, researchers explore the use of Transformer encoder architectures (Vaswani et al. 2017; Tsai et al. 2019a) to achieve alignment between different modalities. Similarly, several studies based on cross-modal attention mechanisms (Goncalves and Busso 2022; Wu et al. 2024; Paraskevopoulos, Georgiou, and Potamianos 2022) provide insights into more effective strategies for aligning multimodal data. Moreover, knowledge is incorporated into the process of integrating multimodal information. KuDA (Feng et al. 2024) leverages affective knowledge to dynamically guide the model in selecting the dominant modality and adjusting the contribution of each modality, while the KEBR (Zhu et al. 2024) injects non-verbal information from videos into the semantic representations of text, thereby enhancing the multimodal representation of textual data. Although most works make progress in the alignment and integration of multimodal information, they ignore the impact of modality noise, thus limiting the performance of the model. Our work dynamically edits the noise in different modality information to improve the performance of MSA.
Multimodal Sentiment Analysis Denoising
Recently, noise in MSA has attracted the attention of researchers. The t-HNE (Li and Li 2025) removes noise from vision and audio through text guidance and attention mechanism. Meta-NA (Zeng, Zhou, and Liu 2022b) learning strategy uses meta-learning to simulate noise tasks and train the model to deal with noise data. JOSFD (Jiang et al. 2024) introduces fuzzy logic into the multimodal fusion and decision-making process to effectively model the subjective and objective ambiguity of emotions. In addition, missing modality is also regarded as a kind of noise in MSA, and thus receives extensive attention(Zeng, Liu, and Zhou 2022; Li et al. 2024). However, research on missing modality mainly focuses on modality completion(Zeng, Zhou, and Liu 2022a) or reducing the interference of missing modality(Shi et al. 2024), and pays insufficient attention to the problem of noise in the original modality. Although modal noise attracts attention, most methods are too coarse in denoising, treating each modality as an independent denoising unit. This approach may lose key information due to excessive denoising, or retain noise due to incomplete processing. In contrast, this study proposes a fine-grained modal noise dynamic editing strategy to effectively remove noise while retaining important information.
MoLAN Framework
@@
As shown in the purple box in the upper part of Figure 2, the MoLAN framework consists of two components: modality-aware blocking and noise dynamic editing. The framework input is the encoder output of the MSA model.
Modality-Aware Blocking
@@
Since the distribution of noise is uneven, we introduce a block-level mechanism to enable fine-grained control over the denoising range. Through blocking, the minimum unit of the denoising operation shifts from the entire modality feature to a feature block, thereby achieving more precise denoising. Considering the differences between modalities, visual information usually presents in regional forms, which is suitable for two-dimensional blocking. In contrast, audio information appears as continuous segments, so one-dimensional blocking is more effective. Feature representation of modalities as $X^{f}\in R^{bs\times m_{f}\times n_{f}},f\in t,a,v$.
@@
where $P_{block}$ is the optimal blocking parameter, it represents the size of the block. $bs$ is batch size. We use two-dimensional blocking as an example to illustrate the blocking process. $(k^{*}_{v},j^{*}_{v})$ is calculated as:
$(k^{*}_{v},j^{*}_{v})=\underset{(k_{v},j_{v})\in\mathcal{D}_{m_{v}}\times\mathcal{D}_{n_{v}}}{\arg\min}\left\|\frac{k_{v}}{\sqrt{m_{v}}}-1\right\|_{2}^{2}+\left\|\frac{j_{v}}{\sqrt{n_{v}}}-1\right\|_{2}^{2}$ (2)
where $\mathcal{D}_{m}$ and $\mathcal{D}_{n}$ represent the sets of factors $k_{v}$ and $j_{v}$ of $m_{v}$ and $n_{v}$, respectively. The $\|\cdot\|_{2}^{2}$ represents the square of the L2 norm. Using the factor closest to the square root as the basis for block division can achieve balanced segmentation: it can avoid information loss or noise residue caused by too large blocks, and semantic loss caused by too small blocks. The modality feature $X^{v}$ is reshaped according to the block parameters, and each sub-block is defined as:
@@
$\forall\;p\in\left[1,\frac{m_{v}}{k^{*}_{v}}\right],\;q\in\left[1,\frac{n_{v}}{j^{*}_{v}}\right]$
where $B^{v}_{pq}\in R^{bs\times k^{*}_{v}\times j^{*}_{v}}$. $p$ and $q$ index the row and column positions of the sub-blocks. See the supplementary material for more detailed reasons for selecting factors for block partitioning based on square root.
Noise Dynamic Editing
@@
MoLAN+ Methodology
As shown in Figure 2, we propose the MoLAN+ method built upon MoLAN framework. The MoLAN+ method consists of three components: the MoLAN framework, noise-suppressed cross attention, and denoise-driven contrastive learning. The detailed introduction of each module can be found in following subsections.
Problem Definition
@@
To ensure a fair comparison, we use the same feature encoder as previous work (Tsai et al. 2019a; Wu et al. 2024; Sun and Tian 2025a). After encoding, we feed the modality features into the MoLAN to obtain the denoised features:
@@
Noise-Suppressed Cross Attention
@@
$\Omega_{\text{Inter-M}}$ $(X^{f_{q}}_{d},X^{f_{kv}}_{d},X^{f_{kv}}_{d},\tilde{M}_{f_{q}})=$ (12)
@@
where $f_{q},f_{kv}\in\{t,a,v\}$. The text attention mask matrix is generated by the encoder. $X^{f_{q}}_{c}$ represents the feature after the attention mechanism.
Denoise-Driven Contrastive Learning
As shown in the light gray below of Figure 2, we introduce a noise-driven contrastive learning loss. We perform contrastive learning on the denoised features and their corresponding original features to encourage the model to learn more discriminative denoised representations. The contrastive learning loss function is as follows:
@@
$-E_{(X^{v},X^{a})\sim\rho(X^{v},X^{a})}\log\left[\frac{\exp\left(\phi(X^{v}_{d},X^{v})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\phi(X^{v}_{d},X^{v}_{j})/\tau\right)}\right]$
$-E_{(X^{v},X^{a})\sim\rho(X^{v},X^{a})}\log\left[\frac{\exp\left(\psi(X^{a}_{d},X^{a})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\psi(X^{a}_{d},X^{a}_{j})/\tau\right)}\right]$
where $\phi$ and $\psi$ denote similarity measurement functions, with cosine similarity adopted in this work. $\tau$ is a temperature parameter that controls the smoothness of the distribution. During training, this module guides the encoder to focus on distinguishing modal noise, thereby improving the overall denoising quality and the effectiveness of multimodal fusion.
Table 1: The performance of the MoLAN framework on the MOSI and MOSEI. The baseline results in the experiment are obtained through replication. Two evaluation metrics, ACC and F1, are adopted, specifically ACC2Has0 / ACC2Non0 and F1Has0 / F1Non0. We perform significance testing on five experimental groups, with p-value of $5.33\times 10^{-5}$, $1.43\times 10^{-6}$, $4.87\times 10^{-10}$, $2.68\times 10^{-6}$, and $2.16\times 10^{-3}$, all of which $<$ 0.05 indicate significant differences.
Sentiment Prediction
By introducing a cross-attention mechanism, information from different modalities is effectively interacted and integrated, enabling the initial fusion of multimodal features. Subsequently, the interacted representations are fed into a self-attention mechanism to further model the deep intra-modal dependencies. Finally, the information from different modalities is consolidated into a unified representation, which is used for the sentiment analysis task.

### v2

@@ Abstract
Multimodal Sentiment Analysis aims to integrate information from various modalities to make complementary predictions. However, it often struggles with irrelevant or misleading visual and auditory information. Most existing approaches treat entire modality as an independent unit for feature enhancement or denoising, which often suppresses redundant noise at the cost of weakening critical information. To address this challenge, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. Specifically, MoLAN performs modality-aware block partitioning by dividing the features of each modality into multiple blocks. Each block is then dynamically assigned a distinct denoising strength based on its noise level and semantic relevance, enabling fine-grained noise suppression while preserving essential multimodal information. Notably, MoLAN is a unified and flexible framework that can be seamlessly integrated into a wide range of multimodal models. Building upon this framework, we further introduce MoLAN+, a new multimodal sentiment analysis approach. Experiments across five models and four datasets demonstrate the broad effectiveness of the MoLAN framework. Extensive evaluations show that MoLAN+ achieves the state-of-the-art performance. The code is publicly available at https://github.com/betterfly123/MoLAN-Framework.
MoLAN: A Unified Modality-Aware Noise Dynamic Editing Framework for Multimodal Sentiment Analysis
Xingle Xu, Yongkang Liu, Dexian Cai, Shi Fengthanks: Corresponding author., Xiaocui Yang, Daling Wang, Yifei Zhang Northeastern University, China, xuxingle@stumail.neu.edu.cn, misonsky@163.com, 2301840@stu.neu.edu.cn {fengshi, yangxiaocui, wangdaling, zhangyifei}@cse.neu.edu.cn
1 Introduction
Multimodal Sentiment Analysis (MSA) aims to integrate information from various modalities to achieve a more comprehensive and accurate understanding of the emotions Zadeh et al. (2018b); Tsai et al. (2019a). MSA holds significant academic value in advancing multimodal learning, and offers broad industrial applications in areas such as human-computer interaction and mental health monitoring Zhu et al. (2025); Singh et al. (2024).
@@ 1 Introduction
Existing MSA methods leverage multimodal synergy to achieve impressive improvements Wu et al. (2024); Li and Liu (2025), but in real-world scenarios, noise interferes with representation learning and leads to performance degradation Li and Li (2025); Liu (2024). To deal with multiple noise patterns, early solutions train individual models from scratch for each noise type Yuan et al. (2021) or design a unified model to perceive noise adaptively Zeng et al. (2022c). However, noise sources differ substantially across modalities, making cross-modality noise transfer prone to failure. Accordingly, researchers design pattern specific denoising methods Yuan et al. (2024), yet they operate at the whole modality. Moreover, noise intensity varies across regions within the modality, so holistic denoising often suppresses noise at the cost of losing essential information.
@@
As shown in Figure 1, the intensity of this noise varies across different regions. Specifically, background smiles that contradict Ross’s emotional state constitute strong noise in the visual modality, while the remaining background mostly represents weak noise. In the audio modality, segments such as laughter that conflict with the sentiment label also form strong noise. The inconsistency of the noise distribution highlights the importance of fine-grained denoising. Therefore, the challenge of this paper is how to perform fine-grained noise dynamic editing on different modalities, so as to remove noise information while retaining information that is beneficial to MSA.
@@
To address these issues, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. To achieve dynamic fine-grained denoising, MoLAN employs a block partitioning strategy that divides each modality into different sub-blocks. In each block, the denoising strength is dynamically computed based on the noise level of the block. This approach allows the model to apply varying degrees of denoising across different blocks, thereby enhancing its ability of selective noise editing while preserving essential information in each modality. In addition, considering the heterogeneity between different modalities Fan et al. (2024); Wei et al. (2023), we design differentiated block partitioning strategies for each modality. Combining our experimental results and following the conclusions of previous studies Li and Li (2025); Zhang et al. (2023); Lin and Hu (2022), we choose text modality as the main basis to calculate the denoising strength. Furthermore, MoLAN is a unified framework that can be flexibly integrated into different MSA models and Multimodal Large Language Models (MLLMs), thereby raising the upper limit of model performance. Based on the MoLAN, we propose MoLAN+, which uses denoised information to update the cross-attention between modalities and guide the model to focus on important information for MSA. MoLAN+ further introduces denoising-driven contrastive learning to encourage the model to generate higher quality features, improving the performance of MSA task. The key contributions are as follows:
@@
  • To address the noise, we propose MoLAN, a unified ModaLity-aware noise dynAmic editiNg framework. It performs modality-aware block partitioning by dividing modality into multiple blocks. Each block is dynamically assigned a distinct denoising strength, enabling fine-grained noise editing. Additionally, MoLAN can be flexibly integrated into various models.
@@
  • We conduct experiments on seven models and four datasets to demonstrate the broad effectiveness of the MoLAN framework. Additionally, extensive evaluations on four benchmark multimodal datasets show that MoLAN+ achieves the state-of-the-art performance.
2 Related Work
2.1 Multimodal Sentiment Analysis (MSA)
MSA enables machines to understand emotions by leveraging visual, audio, and text signals. Early studies mainly adopt fusion methods such as TFN Zadeh et al. (2017) and LMF Liu and Shen (2018) to obtain joint representations. Subsequently, Transformer encoder architectures Vaswani et al. (2017) and cross-modal attention become mainstream. For example, MulT Tsai et al. (2019a) uses cross-modal attention to align and fuse modalities, and related work Zhou et al. (2025); Wu et al. (2024); Guo et al. (2024) further explores more effective alignment strategies. More recently, knowledge is also incorporated. KuDA Feng et al. (2024) leverages affective knowledge to dynamically select the dominant modality and adjust modality contributions, while KEBR Zhu et al. (2024) injects non-verbal information from videos into textual semantics to enhance representations. Despite continuous progress in alignment and fusion, the impact of modality noise is often overlooked, which limits model performance. This work focuses on noise across modalities and performs noise editing to improve MSA.
@@ 2.1 Multimodal Sentiment Analysis (MSA)
Figure 2: An illustration of MoLAN framework and MoLAN+ method. The purple box above represents the MoLAN framework, and the below represents the entire process of the MoLAN+ method. The MoLAN framework shown above provides a detailed description of the MoLAN block presented in MoLAN+ method.
2.2 Multimodal Sentiment Analysis Denoising
Recently, noise in MSA attracts increasing attention. t-HNE Li and Li (2025) removes visual and audio noise via text guidance and attention mechanisms. Meta-NA Zeng et al. (2022c) simulates noise tasks through meta-learning to improve robustness. JOSFD Jiang et al. (2024) introduces fuzzy logic into multimodal fusion and decision-making to model emotion uncertainty. Missing modality is also regarded as a type of noise. EMMR Zeng et al. (2022b) reconstructs semantic features of key missing modalities, TATE Zeng et al. (2022a) uses tags to guide the model to focus on missing information, IASE Shi et al. (2024) aggregates data with a bipartite-graph formulation to reduce the impact of missing modalities, UMDF Li et al. (2024b) learns strong representations via a unified self-distillation mechanism, and Prompts are also used for missing modalities Guo et al. (2024). Unfortunately, existing works adopt too coarse processing granularity in the process of noise removal. This defect may cause excessive denoising and loss of essential information, or may lead to incomplete noise removal. In contrast, our work performs fine-grained noise dynamic editing on modality features, ensuring that essential information is preserved while denoising.
3 MoLAN Framework
@@
As shown in the purple box in the upper part of Figure 2, the MoLAN framework consists of two components: modality-aware block partitioning and noise dynamic editing. The pilot study is provided in Appendix A.
3.1 Modality-Aware Block Partitioning
@@
Since the distribution of noise is uneven, we introduce a block-level mechanism to enable fine-grained control over the denoising range. Through block partitioning, the minimum unit of the denoising operation shifts from the entire modality feature to a feature block, thereby achieving more precise denoising. Considering the differences between modalities, visual information usually presents in regional forms, which is suitable for two-dimensional block partitioning. In contrast, audio information appears as continuous segments, so one-dimensional block partitioning is more effective. Feature representation of modalities as $X^{f}\in R^{bs\times m_{f}\times n_{f}},f\in t,a,v$.
@@
where $P_{block}$ is the optimal block partitioning parameter, it represents the size of the block. $bs$ is batch size. We use two-dimensional block partitioning as an example to illustrate the block partitioning process. $(k^{*}_{v},j^{*}_{v})$ is calculated as:
$(k^{*}_{v},j^{*}_{v})=$ (2)
$\underset{(k_{v},j_{v})\in\mathcal{D}_{m_{v}}\times\mathcal{D}_{n_{v}}}{\arg\min}\left\|\frac{k_{v}}{\sqrt{m_{v}}}-1\right\|_{2}^{2}+\left\|\frac{j_{v}}{\sqrt{n_{v}}}-1\right\|_{2}^{2}$
where $\mathcal{D}_{m}$ and $\mathcal{D}_{n}$ represent the sets of factors $k_{v}$ and $j_{v}$ of $m_{v}$ and $n_{v}$, respectively. The $\|\cdot\|_{2}^{2}$ represents the square of the L2 norm. Using the factor closest to the square root as the basis for block partitioning can achieve balanced segmentation: it can avoid information loss or noise residue caused by too large blocks, and semantic loss caused by too small blocks. The modality feature $X^{v}$ is reshaped according to the block partitioning parameters, and each sub-block is defined as:
@@
$\forall\;p\in\left[1,\frac{m_{v}}{k^{*}_{v}}\right],\;q\in\left[1,\frac{n_{v}}{j^{*}_{v}}\right]\$
where $B^{v}_{pq}\in R^{bs\times k^{*}_{v}\times j^{*}_{v}}$. $p$ and $q$ index the row and column positions of the sub-blocks. We conduct ablation studies to explain the choice of the block partitioning factor.
3.2 Noise Dynamic Editing
@@
4 MoLAN+ Methodology
As shown in Figure 2, we propose the MoLAN+ method built upon MoLAN framework. The MoLAN+ method consists of three components: the MoLAN framework, noise-suppressed cross attention, and denoising-driven contrastive learning. The detailed introduction of each module can be found in following subsections.
4.1 Problem Definition
@@
To ensure a fair comparison, we use the same feature encoder as previous work Tsai et al. (2019a); Wu et al. (2024); Sun and Tian (2025b). After encoding, we feed the modality features into the MoLAN to obtain the denoised features:
@@
4.2 Noise-Suppressed Cross Attention
@@
$\Omega_{\text{Inter-M}}$ $(X^{f_{q}}_{d},X^{f_{kv}}_{d},\tilde{M}_{f_{q}})=$ (12)
@@
where $f_{q},f_{kv}\in\{t,a,v\}$. $f_{q}$ represents the query source modality of the current modality, and $f_{kv}$ represents the modality that provides key/value pairs. The text attention mask matrix is generated by the encoder. $X^{f_{q}}_{c}$ represents the feature after the attention mechanism.
4.3 Denoising-Driven Contrastive Learning
As shown in the light gray below of Figure 2, to enhance the encoder’s capability in distinguishing noise, we introduce a noise-driven contrastive learning loss. We perform contrastive learning between the denoised modal features and their corresponding original features. By minimizing the distance between positive pairs and maximizing the differences with other samples, the model is encouraged to learn more discriminative denoised representations. The contrastive learning loss function can be formulated as follows:
@@ 4.3 Denoising-Driven Contrastive Learning
$-E\log\left[\frac{\exp\left(\phi(X^{v}_{d},X^{v})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\phi(X^{v}_{d},X^{v}_{j})/\tau\right)}\right]$
$-E\log\left[\frac{\exp\left(\psi(X^{a}_{d},X^{a})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\psi(X^{a}_{d},X^{a}_{j})/\tau\right)}\right]$
where $\phi$ and $\psi$ denote similarity measurement functions, with cosine similarity adopted in this work. $\tau$ is a temperature parameter that controls the smoothness of the distribution. During training, this module guides the encoder to focus on distinguishing modal noise, thereby improving the overall denoising quality and the effectiveness of fusion.
4.4 Sentiment Prediction
By introducing a cross-attention mechanism, information from different modalities is effectively interacted and integrated, enabling the initial fusion of multimodal features. Subsequently, the interacted representations are fed into a self-attention mechanism to further model the deep intra-modal dependencies. Finally, the information from different modalities is consolidated into a unified representation, which is used for the sentiment analysis.